# New Construction of Locally Perfect Nonlinear Functions with Application to Sequences Sets with Low Ambiguity Zone

### v2

@@ IV New LAZ sequence sets
The parameters $M$, $D$, $Z_{x}$, $Z_{y}$, $\theta_{\text{max}}$ and $\hat{\theta}_{\text{max}}$ can be seen directly from Theorem 3 and 5 by taking $s=2$ and $R=1$. By Theorem 1, the YZLFLT bounds $\Delta$ and $\hat{\Delta}$ can be derived as follows:
$\left\{\begin{array}[]{ll}\Delta=\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D}{D(MZ_{x}-1)}}=N\sqrt{K}\sqrt{\frac{p-2}{N(p-1)-1}},\\ \hat{\Delta}=\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D-Z_{x}+1}{(MZ_{x}-1)(D+Z_{x}-1)}}=N\sqrt{K}\sqrt{\frac{NK(p-1)-(NK+p-2)}{(NK+p-2)(N(p-1)-1)}},\end{array}\right.$
thus, we have
$\frac{\Delta}{\theta_{\text{max}}}$ $=$ $\frac{N\sqrt{K}\sqrt{\frac{p-2}{N(p-1)-1}}}{K}$
$=$ $\frac{N\sqrt{\frac{p-2}{N(p-1)-1}}}{\sqrt{K}}.$
Observe that if $p\rightarrow\infty$, then $\lim\limits_{p\to\infty}\frac{\theta_{\text{max}}}{\Delta}=\lim\limits_{p\to\infty}\sqrt{\frac{K}{N}}=\sqrt{2}$ since $\frac{K}{N}=\frac{p(p+2)}{\frac{p^{2}-1}{2}}=\frac{2p^{2}+4p}{p^{2}-1}\rightarrow 2$. Furthermore, we have
$\frac{\hat{\Delta}}{\hat{\theta}_{\text{max}}}$ $=$ $\frac{N\sqrt{K}\sqrt{\frac{NK(p-1)-(NK+p-2)}{(NK+p-2)(N(p-1)-1)}}}{K+p-2}$
$=$ $\frac{N\sqrt{K}\sqrt{\frac{NK(p-2)-(p-2)}{NKN(p-1)+N(p-2)(p-1)-(p+NK-2)}}}{K+p-2}.$
Observe that if $p\rightarrow\infty$, then $\lim\limits_{p\to\infty}\frac{\hat{\Delta}}{\hat{\theta}_{\text{max}}}=\lim\limits_{p\to\infty}\frac{N\sqrt{K}}{K}\sqrt{\frac{NK(p-2)}{NKN(p-1)}}=\lim\limits_{p\to\infty}\sqrt{\frac{N}{K}}$ and $\lim\limits_{p\to\infty}\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}=\lim\limits_{p\to\infty}\sqrt{\frac{K}{N}}=\sqrt{2}$ since $\hat{\theta}_{\text{max}}=p(p+2)+p-2\rightarrow K$ and $\frac{K}{N}\rightarrow 2$. The proof of Corollary 2 is completed. ∎
Similarly, we can derive the following corollary, which addresses the case where $K$ is the product of distinct odd primes.
Corollary 3.
Let $K=p_{1}p_{2}\cdots p_{s}$, where $p_{1}<p_{2}<\cdots<p_{s}$ are odd primes, and when $N=E=\textrm{lcm}\{\varphi(p_{i})=p_{i}-1:1\leq i\leq s\}=\frac{(p_{1}-1)(p_{2}-1)\cdots(p_{s}-1)}{2}$. There exist LAZ sequence set with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N=\frac{(p_{1}-1)(p_{2}-1)\cdots(p_{s}-1)}{2}$, $D=NK=\frac{(p_{1}-1)(p_{2}-1)\cdots(p_{s}-1)p_{1}p_{2}\cdots p_{s}}{2}$, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K=p_{1}p_{2}\cdots p_{s}$, $Z_{x}=p_{1}-1$, $\theta_{\text{max}}=K=p_{1}p_{2}\cdots p_{s}$, and $\hat{\theta}_{\text{max}}=K+Z_{x}-1=p_{1}p_{2}\cdots p_{s}+p_{1}-2$. Moreover, $\lim\limits_{D\to\infty}\rho_{\text{LAZ}}=\sqrt{2}$ and $\lim\limits_{D\to\infty}\hat{\rho}_{\text{LAZ}}=\sqrt{2}$, where $\rho_{\text{LAZ}}=\frac{\theta_{\text{max}}}{\Delta}$ and $\hat{\rho}_{\text{LAZ}}=\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}$.
Proof.
The method is the same as in Corollary 1 and 2, and thus it is omitted. ∎
Here, we provide an experimental example demonstrating that values of $N$ can be identified for which the sequence set $\mathcal{S}$ is a LAZ sequence set, achieves asymptotic optimality. Thus, with careful selection of parameters, our sequences are strong candidates for ISAC system.
@@ IV-A Example of Periodic and Aperiodic LAZ sequence set
In this subsection, we present two examples to illustrate our proposed LAZ sequence sets in Theorems 5 and 6.
@@ IV-B Parameters of the Derived Periodic and Aperiodic LAZ Sequence Sets
In Tables II, III, and IV, we present several parameters of the proposed periodic LAZ sequence sets in Theorems 5 and 6, along with their corresponding optimality factor, $\rho_{\text{LAZ}}.$ Furthermore, we provide parameters of the proposed aperiodic LAZ sequence sets in Theorems 5 and 6, along with their corresponding optimality factor, $\hat{\rho}_{\text{LAZ}}$, in Tables V and VI, respectively. It is obvious from these tables that the optimality factors ${\rho}_{\mathrm{LAZ}}$ and $\hat{\rho}_{\text{LAZ}}$ of the constructed LAZ sequence sets are asymptotically approaching $1$. Thus, with careful selection of parameters, our sequences are strong candidates for ISAC system.
@@
$N$ $NK$ $\Pi$ magnitude $\theta_{\max}$ $\rho_{\mathrm{LAZ}}$
@@
$N$ $NK$ $\Pi$ magnitude $\theta_{\max}$ $\rho_{\mathrm{LAZ}}$
@@
$N$ $NK_{1}$ $\Pi$ magnitude $\theta_{\max}$ $\rho_{\mathrm{LAZ}}$
@@
$N$ $NK$ $\Pi$ magnitude $\theta_{\max}$ $\rho_{\mathrm{LAZ}}$
@@
$N$ $NK_{1}$ $\Pi$ magnitude $\hat{\theta}_{\max}$ $\hat{\rho}_{\mathrm{LAZ}}$
@@ V Concluding Remarks
In this paper, we introduce some constructions of Locally Perfect Nonlinear Functions (LPNFs) with new parameters. By leveraging their structural features, we further propose a series of LAZ sequence sets with more flexible parameters. Notably, one class of these sequence sets achieves both periodic and aperiodic asymptotic optimality.

### v3

@@ IV Proposed LAZ sequence sets
By (7) and (9), we obtain the YZLFLT bounds $\Delta$ and $\hat{\Delta}$:
$\left\{\begin{array}[]{ll}\Delta=\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D}{D(MZ_{x}-1)}}=\sqrt{\frac{D(1-\frac{D}{MZ_{x}Z_{y}})}{1-\frac{1}{MZ_{x}Z_{y}}}},\\ \hat{\Delta}=\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D-Z_{x}+1}{(MZ_{x}-1)(D+Z_{x}-1)}}=\sqrt{\frac{D(1-\frac{D}{MZ_{x}Z_{y}}-\frac{1}{MZ_{y}}+\frac{1}{MZ_{x}Z_{y}})}{(1-\frac{1}{MZ_{x}})(1+\frac{Z_{x}}{D}-\frac{1}{D})}}.\end{array}\right$. (13)
(i)  We use $\Delta$ in (13) with $M=N=p^{e-1}(p-1)$, $D=NK_{1}$, $Z_{y}=K_{1}-K+2$, $Z_{x}=p-1$ to get
$\Delta$ $=$ $\sqrt{\frac{D(1-\frac{K_{1}}{(p-1)(K_{1}-K+2)})}{1-\frac{1}{p^{e-1}(p-1)^{2}(K_{1}-K+2)}}}.$ (14)
Moreover $\theta_{\text{max}}=K_{1}$, $\lim\limits_{p\to\infty}\sqrt{\frac{K_{1}}{N}}=1$, and $\lim\limits_{p\to\infty}\frac{K_{1}}{(p-1)(K_{1}-K+2)}=0$, it follows that
$\lim\limits_{p\to\infty}\frac{\theta_{\text{max}}}{\Delta}=\lim\limits_{p\to\infty}\sqrt{\frac{K_{1}^{2}}{D}}=\lim\limits_{p\to\infty}\sqrt{\frac{K_{1}}{N}}=1.$
(ii)  We use $\hat{\Delta}$ in (13) with $M=N=p^{e-1}(p-1)$, $D=NK_{1}$, $Z_{y}=K_{1}-K+2$, $Z_{x}=p-1$ to get
$\hat{\Delta}$ $=$ $\sqrt{\frac{D(1-\frac{K_{1}}{(p-1)(K_{1}-K+2)}-\frac{1}{p^{e-1}(p-1)(K_{1}-K+2)}+\frac{1}{p^{e-1}(p-1)^{2}(K_{1}-K+2)})}{(1-\frac{1}{p^{e-1}(p-1)^{2}})(1+\frac{1}{p^{e-1}K_{1}}-\frac{1}{p^{e-1}(p-1)K_{1}})}}.$ (15)
We observe that if $e\geq 2$ and $p\rightarrow\infty$, combining $\hat{\theta}_{\text{max}}=K_{1}+p-2$, we have
$\lim\limits_{p\to\infty}\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}=\lim\limits_{p\to\infty}\sqrt{\frac{(K_{1})^{2}+2K_{1}(p-2)+(p-2)^{2}}{NK_{1}}}=\lim\limits_{p\to\infty}\sqrt{\frac{K_{1}}{N}}=1$
since $\lim\limits_{p\to\infty}\sqrt{\frac{K_{1}}{N}}=1$ and $\lim\limits_{p\to\infty}\frac{K_{1}}{(p-1)(K_{1}-K+2)}=0$. The proof of Theorem 8 is complete. ∎
For the $(N,NK,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$-LAZ sequence set constructed in [18], the value of the parameter $N$ is only odd. But for the construction in Theorem 1 (i), the value of parameter $N$, as a factor of $\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}$, can be chosen to be even. Thus the LAZ sequence set in Theorem 3 has more general and flexible parameters. Moreover, we also present two other new classes of constructions of LPNFs in Theorem 1 (ii) and Theorem 2, which are used to design more LAZ sequence sets with flexible parameters.
@@ IV-A Example of Periodic and Aperiodic LAZ sequence set
In this subsection, we present two examples to illustrate our proposed LAZ sequence sets in Theorems 3, 5 and 7.
@@ IV-B Parameters of the Derived Periodic and Aperiodic LAZ Sequence Sets
In Tables II, III, and IV, we display several parameters of the proposed periodic LAZ sequence sets in Theorems 3 and 7, along with their corresponding optimality factor, $\rho_{\text{LAZ}}.$ Furthermore, we provide parameters of the proposed aperiodic LAZ sequence sets in Theorems 5 and 7, along with their corresponding optimality factor, $\hat{\rho}_{\text{LAZ}}$, in Tables V and VI, respectively. It is obvious from these tables that the optimality factors ${\rho}_{\mathrm{LAZ}}$ and $\hat{\rho}_{\text{LAZ}}$ of the constructed LAZ sequence sets are asymptotically approaching $1$. Thus, with careful selection of parameters, our sequences are strong candidates for ISAC system.
@@
NN NKNK $\Pi$ magnitude $\theta_{\max}$ $\rho_{\mathrm{LAZ}}$
@@
NN NKNK $\Pi$ magnitude $\theta_{\max}$ $\rho_{\mathrm{LAZ}}$
@@
NN $NK_{1}$ $\Pi$ magnitude $\theta_{\max}$ $\rho_{\mathrm{LAZ}}$
@@
NN NKNK $\Pi$ magnitude $\theta_{\max}$ $\rho_{\mathrm{LAZ}}$
@@
NN $NK_{1}$ $\Pi$ magnitude $\hat{\theta}_{\max}$ $\hat{\rho}_{\mathrm{LAZ}}$
@@ V Concluding Remarks
We investigated three constructions of LPNFs with new parameters. By leveraging their structural features, we further proposed a series of periodic and aperiodic LAZ sequence sets with more flexible parameters. In comparison with prior works, we provided more new LAZ sequence sets, owing to reduced constraints on parameters. Significantly, two classes of these newly developed periodic and aperiodic LAZ sequence sets are asymptotically optimal based on YZFLLT bounds. Moreover, another class of periodic LAZ sequence sets is asymptotically optimal based on YZFLLT bounds.